# Scalable Dynamic Origin-Destination Demand Estimation Enhanced by High-Resolution Satellite Imagery Data

### v1

@@ 3 Model
This section introduces the DODE framework leveraging high-resolution satellite images as an additional data source. Figure 1 demonstrates overall framework consisting three parts: (1) multi-source data collection, including satellite imagery, road network and tabular data from localized ground-based sensors. (2) a computer vision pipeline to extract traffic information from satellite images and (3) DODE model incorporating all available data sources. In the following subsections, we first introduce the computer vision pipeline for density extraction. Next we formulate the DODE problem using traffic information from local ground-based detectors and satellite images with an introduction of a general dynamic network loading model. Furthermore, we present the framework on a computational graph to solve more efficiently using forward-backward algorithms. Lastly, we discuss data availability issues and corresponding aggregation methods.
Figure 1: Model Overview
@@ 3.1.1 Vehicle detection
Vehicle detection in satellite imagery is a well-studied problem in computer vision, and various pre-trained models have been open-sourced [Nachmany and Alemohammad, 2019]. Liu et al. [2024] compared the performance of two model architectures provided in the Python deep learning library Raster Vision [Azavea/Element 84, Robert Cheetham, n.d.]. The first architecture is a Single Shot Multibox Detector (SSD) model [Liu et al., 2016], which uses ”VGG” as backbone [Simonyan and Zisserman, 2014]. The second architecture is a Faster R-CNN (FRCNN) model [Ren et al., 2017] that uses Resnet50 as backbone [He et al., 2016]. Liu et al. [2024] found that FRCNN gives the best results in terms of precision, recall and F1-score, and this study adopts the same Faster R-CNN architecture but extends its application to a multi-class classification setting, enabling the model to distinguish between cars and trucks.
The object detection model was trained using high-resolution satellite imagery from the xView dataset [Lam et al., 2018], which contains 744 annotated images at a 30-centimeter resolution. The dataset includes 60 object categories, 26 of which correspond to various types of vehicles. These 60 child categories are grouped into 7 parent classes. When labelers are unable to assign an object to a specific child category confidently, they fall back to the more general parent category [Lam et al., 2018]. Instead of classifying each object into its original xView category, the model was designed to perform a coarser multi-class classification task by categorizing detected objects as either car, truck, other vehicle, or non-vehicle. The car class includes the parent class Passenger Vehicle and the child class Small Car. The truck class includes the parent class Truck and nine truck-related child categories, such as Cargo Truck, Truck Tractor, and Haul Truck.
@@ 3.1.2 Map matching and density calculation
Raster Vision provides built-in functionality to extract the geospatial coordinates of the bounding boxes associated with cars and trucks in the satellite image. Using standard geospatial analysis tools, the centroids of these bounding boxes were computed and subsequently matched to the nearest links in the street network. To avoid assigning vehicles parking far from road segments or on segments not included in the network, only centroid points of vehicles falling within a predefined buffer around each road centerline were considered eligible for map matching.
After cars and trucks are matched to the links of the street network, traffic density for each link can be calculated as the ratio between the number of matched vehicles and the corresponding link length. This computation is performed separately for cars and trucks, as the DODE framework is capable of incorporating multi-class traffic density data. Because the link lengths are exogenous information, we use the total detected and matched vehicle number by class to represent link density without loss of generality.
@@ 3.2 Dynamic Origin-destination Demand Estimation (DODE)
This section formulates the DODE problem that integrates multi-source mobility data, including link-level traffic count, travel time and additional traffic density extracted from satellite images. We begin by describing the approach to modeling link-level densities for different vehicle classes in mesoscopic DNL model. Next, we formulate the DODE problem and present it on a computational graph and derive the forward and backward processes. Last, we discuss potential issues that may arise when deploying this framework in real-world scenarios with various data availabilities and the corresponding aggregation approaches.
@@ 3.2.1 Preliminary
Let $G=(N,A)$ represent a general road network with a node set $N$ and a link set $A$. A subset of links, denoted as $A_{p}\subset A$, represents links allowing for on-street parking. The assignment horizon is defined as $T_{d}=[1,2,...,T]$. $R\subset N$ and $S\subset N$ are sets of origin and destination nodes respectively. Between an origin-destination (OD) pair $rs$ where $r,s\in R,S$, $q^{rs}_{i,t}$ represents the total demand of vehicle class $i$ departing at time $t\in T_{d}$ from $r$ and heading to $s$. We consider two vehicle classes in this study: cars denoted by $\mathcal{C}$ and trucks denoted by $\mathcal{T}$. $P^{rs}_{i}$ is the set of paths available for vehicle class $i$ to choose between the OD pair $rs$, and $f^{rs}_{k,i,t}$ represents the path flow (number of trips) of vehicle class $i$ departing at time $t$ and choosing path $k\in P^{rs}_{i}$. The corresponding path travel cost is denoted as $c^{rs}_{i,k,t}$. $p^{rs}_{k,i,t}$ represents route choice proportion of choosing path $k$ for vehicle class $i$ departing from $r$ at time $t$. Therefore, path flow $f^{rs}_{k,i,t}$ can be obtained by multiplying the route choice proportion $p^{rs}_{k,i,t}$ with OD demand $q^{rs}_{i,t}$.
@@
$\bm{f}_{i}=\bm{p}_{i}\bm{q}_{i},\ \forall i\in C$ (2)
The route choice portions $p^{rs}_{k,i,t}$ can be determined using various route choice function, such as dynamic user equilibrium [Pi et al., 2019, Liu and Qian, 2024] and stochastic user equilibrium based on Logit model [Guarda et al., 2024]. The inputs depend on the specific formulation and here we use path and link level travel costs, denoted by $c^{rs}_{k,i,t}$ and $h^{t}_{a,i}$, without loss of generality.
@@
$\bm{p}_{i}=\Psi(\{\bm{c}_{i}\}_{i},\{\bm{{h}}_{i}\}_{i}),\ \forall i\in C$ (4)
In this study, the agent-based mesoscopic dynamic network loading (DNL) model, MAC-POSTS [Pi et al., 2018, Ma et al., 2020], is used to simulate network-wide time-varying traffic conditions, including traffic count, travel time and density. In the DNL, cumulative curves are installed at the head and tail of each observed link to extract time-varying traffic conditions.
For links with traffic flow and/or travel time observations, we use two cumulative curves recording cumulative numbers of arriving and departing through vehicles respectively on link $a$, denoted by $\mathcal{A}^{m}_{a,i}$ and $\mathcal{D}^{m}_{a,i}$ respectively. Note that these vehicles do not stop (i.e., on-street parking) on this link and First-In-First-Out (FIFO) is enforced, hence they can be used to extract link travel times. For links with density observations, especially links in $A_{p}$ with on-street parking spaces, we install two additional cumulative curves at the head and tail of link $a$, denoted by $\mathcal{A}^{p}_{a,i}$ and $\mathcal{D}^{p}_{a,i}$ respectively, to record in/outflow for parking vehicles. Due to stochastic parking durations, FIFO is not enforced in this pair of cumulative curves and they are not used to extract travel time. Figure 2 demonstrates the setting of two pairs of cumulative curves at observed links with on-street parking spaces. To efficiently extract DAR matrices, these cumulative curves are tree-based [Ma et al., 2020, Liu and Qian, 2024].
@@
We incorporate curbside parking dynamics into the DNL model using the approach proposed in Liu and Qian [2024]. Truncated fundamental diagrams with respect to curb occupancies are used to govern link dynamics, which can be generalized to any link dynamics model, such as link queue or link transmission models. For instance, for link queue model, we can use link-level parking queues (instead of curb cells in CTM) to store parking vehicles, with the same curb parking dynamics to track curb availability and departure times.
Using these two pairs of cumulative curves, traffic count of class $i$ during time interval $t$ is modeled as the total arrival vehicle number of both parking and through vehicles during time interval $t$, shown as
@@
Link travel time at time $t$ is modeled as the traversing time for vehicles arriving at time $t$ on link $a$, extracted from cumulative curves of through vehicles, which reads
@@ 3.2.2 Modeling traffic density
Traffic density is a measure of the number of vehicles per unit length of road. Because road length are exogenous, we use the total number of vehicles remaining on link $a$ at timestamp $\bar{t}$, denoted by $R_{a,i}(\bar{t})$ to represent the link density, which can be modeled as summation of remaining through and parking vehicle numbers at time $t$, each of them is the difference between cumulative arrival vehicle number and cumulative departure vehicle number before timestamp $t$. As our DNL model is discrete in time, $R_{a,i}(\bar{t})$ (the upper bound) is used to represent the density for time interval $t$.
@@
To obtain more stable and smooth density estimation, we average the densities across several neighboring time intervals around time ${t}$, denoted by $[{t}-\delta,{t}+\delta]$, the average density can be computed by
@@
To make simulation-based DODE problem tractable, we linearize the DNL process using Dynamic Assignment Ratio (DAR) matrix [Ma et al., 2020], which mapping time-varying path flows to time-varying link flows. Because both arrival and departure link flows are used in estimating remaining vehicle numbers, and both parking and non-parking vehicles are accounted in link inflows, four separate DAR matrices are used to record corresponding mapping relationships. The traffic flow, represented as link inflow in our DNL, during time $t_{2}$ has the following relationship with path flows
@@
where $\rho_{\text{arr},k,a,i}^{rs,n/p}(t_{1},t_{2})$ represents the time-dependent incidence matrix, which is the proportion of the path flow $f^{rs}_{k,i,t_{1}}$ contributing to the link arrival flow $x^{t_{2},n/p}_{\text{arr},a,i}$. Here we use $t_{1}$ and $t_{2}$ to differentiate time index for path and link flows respectively. The corresponding vectorized form can be written as
@@
Similarly, the departure flow of link $a$ at time $t_{2}$ for class $i$ can be modeled using departure DAR matrix as
@@
Then the cumulative arrival and departure flows on link $a$ at time ${t}$ for vehicle class $i$ can be computed as the summation of all arrival or departure flows before time ${t}$, that are arrival or departure flows of intervals from $1$ to $t$,
@@
where $\bm{H}_{i}$ is used to aggregate total arrival and departure flow from time interval $1$ to $t_{2}$ in Equation 14.
To summarize, the DNL model, denoted by $\Lambda$, uses time-varying path flows as inputs and gives out link level traffic conditions and DAR matrices mapping recording the mapping relations between path flows and link flows, shown in Equation (16).
@@ 3.2.3 DODE formulation
The DODE can be formulated as a mathematical programming problem shown in Equation 18. The decision variable is OD demand $\{\bm{q}_{i}\}_{i}$. The objective function is minimizing the differences between modeled and observed traffic conditions including link traffic count, link travel time and link density. The constraints include route choice, DNL, path-demand relationship and non-negativity constraint of demand.
@@
$\bm{q}_{i}\geq 0,\ \forall i\in C$
where $\bm{x}^{o}$, $\bm{h}^{o}$ and $\bm{k}^{o}\odot\bm{l}$ are observed count, travel time and remaining vehicles (density multiplied by link length). $\bm{L}_{i}$, $\bm{M}_{i}$ and $\bm{I}_{i}$ are aggregation matrices used to map simulated conditions to available observations, depending on data availability which will be discussed in section 3.4.
@@ 3.3 Computational graph-based solution algorithm
The formulation can be presented on a computational graph and solved by forward-backward algorithms efficiently. In the forward pass shown in Figure 3, a dynamic traffic assignment problem is solved using behavior function $\Psi$ and DNL model $\Lambda$, given fixed time-varying multi-class demand $\{\bm{q}_{i}\}_{i}$ (initialized demand or estimated demand from previous epoch). Route choices $\{\bm{p}_{i}\}_{i}$, link-level traffic states and DAR matrices $\bm{\rho}^{m}_{a,i},\bm{\rho}^{p}_{a,i},\bm{\rho}^{m}_{d,i},\bm{\rho}^{p}_{d,i}$ are obtained from Equations (17), (4) and (2). To improve computation efficiency in DNL on large-scale networks, we can use hybrid route choice models which is embedded in the DNL model to retrieve link traffic measurements and DAR matrices in one DNL run. In the hybrid route choice model, a fraction of travelers follow pre-trip route choice model and the other travelers adaptively explore time-varying shortest paths following en-route route choice model [Qian and Zhang, 2013, Liu and Qian, 2024].
In the backward pass, shown in Figure 3, the gradient of the loss function with respect to demand is derived using the Chain Rule in Equation 19. The gradient of objective function with respect to decision variable $\bm{q}_{i}$ is shown in Equation 20. Any gradient-based solution algorithm can be used to solve the DODE based on this gradient on the computational graph.
@@
3.4 Data availability and aggregation
This section addresses issues related to data availability and aggregation methods. The DNL model simulates network-wide traffic conditions in various resolutions of interest (in this study, we estimate traffic conditions of each link every 15 minutes), however, observed data may not be collected at the same frequency. Moreover, data quality and pre-processing error cannot provide precise spatio-temporal features. To address this, we use aggregation matrices to mapping modeled conditions to observations and aggregation can occur in both spatial and temporal domains.. Temporal aggregation refers to aggregating traffic conditions into larger time intervals. For example, summing up four 15min traffic count to match hourly traffic flow observation [Ma et al., 2020, Liu and Qian, 2024].
Spatial aggregation primarily arises from limitations in satellite image quality and computer vision method’s performance. We discuss three potential scenarios which require spatial aggregation. The first scenario is about lane-level vehicle detection. when CV pipeline does not have precise road segmentation feature, it cannot conduct lane-level vehicle detection, failing to differentiate between parking and through vehicles. In such cases, observations provide link density for all vehicles on the link and we aggregate parking and through vehicles together to model density, which is already described in Equation (7) and (8). In this study, we integrate the aggregation into density estimation in the DNL model. Another general way is estimating separate parking density and through traffic density in the DNL and using matrix $\bm{I}_{i}$ to aggregate them. Another reason we aggregating them in DNL process is that one through vehicle captured by satellite image only represent the status at the timestamp of the image and we are not sure if this vehicle will perform on-street parking after this timestamp or just left a parking spot before this timestamp. Aggregating through and parking vehicles into total density can overcome this information gap. The second scenario is about vehicle moving directions. Similar to scenario 1, if CV pipeline cannot perform road segmentation and detect the vehicle orientations, detected vehicles cannot be correctly matched to directed road segments, potentially leading to flow inconsistency on successive links. To address this, we aggregate two (even more) moving directions together to match regional total density. Another scenario is about vehicle class detection. When the CV pipeline fails to detect vehicle classes with high precision, we can simply aggregate density by vehicle classes.

### v2

@@ 3 Model
This section introduces the DODE framework leveraging high-resolution satellite imagery as an additional data source. Figure 1 demonstrates the overall framework, consisting of three parts: (1) multi-source data collection, including satellite imagery, road network, and tabular data from localized ground-based sensors. (2) a computer vision pipeline to extract traffic information from satellite imagery and (3) DODE model incorporating all available data sources. In the following subsections, we first introduce the computer vision pipeline for density extraction. Next, we formulate the DODE problem using traffic information from local ground-based detectors and satellite images, with an introduction to a general dynamic network loading model that explicitly models parking and on-road vehicles. Furthermore, we present the framework on a computational graph to solve DODE more efficiently using forward-backward algorithms. Lastly, we discuss data availability issues and corresponding aggregation methods.
@@ 3.1.1 Vehicle detection
Vehicle detection in satellite imagery is a well-studied problem in computer vision, and various pre-trained models are available as open-source implementations [Nachmany and Alemohammad, 2019]. Liu et al. [2024] compared the performance of two model architectures provided in the library Raster Vision [Azavea/Element 84, Robert Cheetham, ]. The first architecture is a Single Shot Multibox Detector (SSD) model [Liu et al., 2016], which uses VGG as the backbone [Simonyan and Zisserman, 2014] and the second one is a Faster R-CNN (FRCNN) model [Ren et al., 2017] that uses ResNet-50 as the backbone [He et al., 2016]. Liu et al. [2024] found that FRCNN achieves better results in terms of precision, recall, and F1-score, so this study adopts the same architecture but extends its application to a multi-class classification setting, enabling the model to distinguish between cars and trucks.
The object detection model was trained using high-resolution satellite imagery from the xView dataset [Lam et al., 2018], which contains 744 annotated images at a 30-centimeter resolution. The dataset includes 60 object categories, 26 of which correspond to various types of vehicles. These 60 child categories are grouped into 7 parent classes. When labelers are unable to assign an object to a specific child category confidently, they assign it to the more general parent category [Lam et al., 2018]. Instead of classifying each object into its original xView category, the model was designed to perform a coarser multi-class classification task by categorizing detected objects as either car, truck, other vehicle, or non-vehicle. The car class includes the parent class Passenger Vehicle and the child class Small Car. The truck class includes the parent class Truck and nine truck-related child categories, such as Cargo Truck, Truck Tractor, and Haul Truck.
@@ 3.1.2 Map matching and density calculation
Raster Vision provides built-in functionality to extract the geospatial coordinates of the bounding boxes associated with cars and trucks in the satellite imagery. Using standard geospatial analysis tools, the centroids of these bounding boxes were computed and subsequently matched to the nearest links in the street network. To avoid assigning vehicles parked far from road segments or on segments not included in the network, only centroid points of vehicles falling within a predefined buffer around each road centerline were considered eligible for map matching.
After cars and trucks are matched to the links of the street network, traffic density for each link can be calculated as the ratio of the number of matched vehicles to the corresponding link length. This computation is performed separately for cars and trucks, as the DODE framework is capable of incorporating multi-class traffic density data. Since the link lengths are exogenous information, we use the total detected and matched vehicle counts by class to represent link density without loss of generality.
@@ 3.2 Dynamic Origin-destination Demand Estimation (DODE)
This section formulates the DODE problem that integrates multi-source mobility data, including link-level traffic counts, travel times, and additional traffic density extracted from satellite images. We begin by describing the approach to modeling link-level densities for different vehicle classes in a mesoscopic DNL model. Next, we formulate the DODE problem and present it on a computational graph, and we derive the forward and backward processes. Lastly, we discuss potential issues that may arise when deploying this framework in real-world scenarios with varying data availability and aggregation approaches.
@@ 3.2.1 Preliminary
Let $G=(N,A)$ represent a general road network with a node set $N$ and a link set $A$. A subset of links, denoted as $A_{p}\subset A$, represents links allowing on-street parking. The assignment horizon is defined as $T_{d}=[1,2,...,T]$. $R\subset N$ and $S\subset N$ are sets of origin and destination nodes, respectively. Between an origin-destination (OD) pair $rs$ where $r,s\in R,S$, $q^{rs}_{i,t}$ represents the total demand of vehicle class $i$ departing at time $t\in T_{d}$ from $r$ and heading to $s$. We consider two vehicle classes in this study: cars denoted by $\mathcal{C}$ and trucks denoted by $\mathcal{T}$. $P^{rs}_{i}$ is the set of paths available for vehicle class $i$ to choose for the OD pair $rs$, and $f^{rs}_{k,i,t}$ represents the path flow (number of trips) of vehicle class $i$ departing at time $t$ and choosing path $k\in P^{rs}_{i}$. The corresponding path travel cost is denoted as $c^{rs}_{i,k,t}$. $p^{rs}_{k,i,t}$ represents the route choice proportion for path $k$ for vehicle class $i$ departing from $r$ at time $t$. Therefore, path flow $f^{rs}_{k,i,t}$ can be obtained by multiplying the route choice proportion $p^{rs}_{k,i,t}$ with OD demand $q^{rs}_{i,t}$.
@@
$\bm{f}_{i}=\bm{p}_{i}\bm{q}_{i},\ \forall i\in\mathcal{C}$ (2)
The route choice proportions $p^{rs}_{k,i,t}$ can be determined using various route choice functions, such as dynamic user equilibrium [Pi et al., 2019, Liu and Qian, 2024] and stochastic user equilibrium based on the Logit model [Guarda et al., 2024]. The inputs depend on the specific formulation and here we use path and link level travel costs, denoted by $c^{rs}_{k,i,t}$ and $h^{t}_{a,i}$, without loss of generality.
@@
$\bm{p}_{i}=\Psi(\{\bm{c}_{i}\}_{i},\{\bm{{h}}_{i}\}_{i}),\ \forall i\in\mathcal{C}$ (4)
In this study, the agent-based mesoscopic dynamic network loading (DNL) model, MAC-POSTS [Pi et al., 2018, Ma et al., 2020], is used to simulate network-wide time-varying traffic conditions, including traffic counts, travel times and density. In the DNL, cumulative curves are installed at the head and tail of each observed link to extract time-varying traffic conditions.
For links with traffic flow and/or travel time observations, we use two cumulative curves recording cumulative numbers of arriving and departing through vehicles , respectively on link $a$, denoted by $\mathcal{A}^{m}_{a,i}$ and $\mathcal{D}^{m}_{a,i}$ respectively. For links with density observations, especially links in $A_{p}$ with on-street parking spaces, we install two additional cumulative curves at the head and tail of link $a$, denoted by $\mathcal{A}^{p}_{a,i}$ and $\mathcal{D}^{p}_{a,i}$ respectively, to record inflows and outflows for parking vehicles.
Figure 2 demonstrates the setting of two pairs of cumulative curves at observed links with on-street parking spaces. To efficiently extract DAR matrices, these cumulative curves are tree-based [Ma et al., 2020, Liu and Qian, 2024].
@@
We incorporate curbside parking dynamics into the DNL model using the approach proposed in Liu and Qian [2024]. Truncated fundamental diagrams with respect to curb occupancies are used to govern link dynamics, which can be generalized to any link dynamics model, such as link queue or link transmission models. For instance, for a link queue model, we can use link-level parking queues (instead of curb cells in CTM) to store parking vehicles, with the same curb parking dynamics to track curb availability and departure times. Because the DNL model runs with its queue-based link and node dynamics, First-In-First-Out (FIFO) is explicitly enforced. It uses separate cumulative curves for parking and through vehicles, ensuring that travel times are extracted from FIFO-preserving cumulative curves of through vehicles without the impact of parking-related non-FIFO behaviors. Using these two pairs of cumulative curves, the traffic count of class $i$ during time interval $t$ is modeled as the total arrival count of both parking and through vehicles during time interval $t$, shown as
@@
Link travel time at time $t$ is modeled as the traversal time for vehicles arriving at time $t$ on link $a$, extracted from cumulative curves of through vehicles, which reads
@@ 3.2.2 Modeling traffic density
Traffic density is a measure of the number of vehicles per unit length of road. Because road length is exogenous, we use the total number of vehicles remaining on link $a$ at timestamp $\bar{t}$, denoted by $R_{a,i}(\bar{t})$ to represent the link density, which can be modeled as the summation of remaining through and parking vehicle numbers at time $t$, each of them is the difference between the cumulative arrival vehicle number and the cumulative departure vehicle number before timestamp $t$. As our DNL model is discrete in time, $R_{a,i}(\bar{t})$ (the upper bound) is used to represent the density for time interval $t$.
@@
To obtain more stable and smooth density estimation, we average the densities across several neighboring time intervals around time t{t}, denoted by ${t}-\delta,{t}+\delta]$. The average density can be computed by
@@
To make the simulation-based DODE problem tractable, we linearize the DNL process using Dynamic Assignment Ratio (DAR) matrices to express the mapping between path flows and link flows in a differentiable form for the computational-graph implementation. The estimated density computation is also built upon DAR matrices. Because both arrival and departure link flows are used in estimating remaining vehicle numbers, and both parking and non-parking vehicles are accounted for link inflows, four separate DAR matrices are used to record the corresponding mapping relationships. The traffic flow, represented as link inflow in our DNL, during time $t_{2}$ has the following relationship with path flows
@@
where $\rho_{\text{arr},k,a,i}^{rs,m/p}(t_{1},t_{2})$ represents the time-dependent incidence matrix, which is the proportion of the path flow $f^{rs}_{k,i,t_{1}}$ contributing to the link arrival flow $x^{t_{2},n/p}_{\text{arr},a,i}$. Here we use $t_{1}$ and $t_{2}$ to differentiate the time index for path and link flows respectively. The corresponding vectorized form can be written as
@@
Similarly, the departure flow of link $a$ at time $t_{2}$ for class $i$ can be modeled using the departure DAR matrix as
@@
Then the cumulative arrival and departure flows on link $a$ at time t{t} for vehicle class $i$ can be computed as the summation of all arrival or departure flows before time t{t}, that are arrival or departure flows of intervals from $1$ to $t$,
@@
where $\bm{H}_{i}$ is used to aggregate total arrival and departure flows from time interval $1$ to $t_{2}$ in Equation 14.
To summarize, the DNL model, denoted by $\Lambda$, uses time-varying path flows as inputs and outputs link-level traffic conditions and DAR matrices that record the mapping relationships between path flows and link flows, shown in Equation (16).
@@ 3.2.3 DODE formulation
The DODE can be formulated as a mathematical programming problem shown in Equation 18. The decision variables are OD demand $\{\bm{q}_{i}\}_{i}$. The objective function minimizes the differences between modeled and observed traffic conditions, including link traffic counts, link travel times, and link density. The constraints include route choice, DNL, the path-demand relationship and non-negativity constraint of demand.
@@
$\bm{q}_{i}\geq 0,\ \forall i\in\mathcal{C}$
where $\bm{x}^{o}$, $\bm{h}^{o}$ and $\bm{k}^{o}\odot\bm{l}$ are observed counts, travel times, and remaining vehicles (density multiplied by link length). $\bm{L}_{i}$, $\bm{M}_{i}$ and $\bm{I}_{i}$ are aggregation matrices used to map simulated conditions to available observations, depending on data availability which will be discussed in section 3.4.
@@ 3.3 Computational graph-based solution algorithm
The formulation can be presented on a computational graph and solved by forward-backward algorithms efficiently. In the forward pass shown in Figure 3, a dynamic traffic assignment problem is solved using the behavior function $\Psi$ and DNL model $\Lambda$, given fixed time-varying multi-class demand $\{\bm{q}_{i}\}_{i}$ (initialized demand or estimated demand from the previous epoch). Route choices $\{\bm{p}_{i}\}_{i}$, link-level traffic states and DAR matrices $\bm{\rho}^{m}_{a,i},\bm{\rho}^{p}_{a,i},\bm{\rho}^{m}_{d,i},\bm{\rho}^{p}_{d,i}$ are obtained from Equations (17), (4) and (2). To improve computation efficiency in DNL on large-scale networks, we can use the hybrid route choice models which is embedded in the DNL model to retrieve link traffic measurements and DAR matrices in one DNL run. In the hybrid route choice model, a fraction of travelers follow a pre-trip route choice model and the other travelers adaptively explore time-varying shortest paths following an en-route route choice model [Qian and Zhang, 2013, Liu and Qian, 2024].
In the backward pass, shown in Figure 3, the gradient of the loss function with respect to demand is derived using the Chain Rule in Equation 19. The gradient of the objective function with respect to the decision variable $\bm{q}_{i}$ is shown in Equation 20. Any gradient-based solution algorithm can be used to solve the DODE based on this gradient on the computational graph.
@@
3.4 Discussions
This section first addresses issues related to data availability and aggregation methods. The DNL model simulates network-wide traffic conditions at various spatio-temporal resolutions of interest, but observed data may not be collected at the same frequency. Moreover, data quality and pre-processing errors may not yield precise spatio-temporal observations. To address this, we use aggregation matrices to map modeled conditions to observations, and aggregation can occur in both spatial and temporal domains. Temporal aggregation refers to aggregating traffic conditions into larger time intervals. For example, summing four 15-min traffic counts to match hourly traffic flow observation [Ma et al., 2020, Liu and Qian, 2024]. In this study, satellite imagery provides snapshot-like density observations with a known capture time window, and each density snapshot is assigned to the corresponding 15-min interval in the DNL model. For consistency, ground-sensor observations are also aggregated to the same 15-minute resolution. We use spatio-temporal aggregation operators to map simulated link states to the exact time intervals where measurements exist. This allows the estimation objective to incorporate heterogeneous data streams in a unified way.
Spatial aggregation primarily arises from limitations in satellite image quality and the computer vision pipeline’s performance. We discuss three potential scenarios that require spatial aggregation. The first scenario is about lane-level vehicle detection. When CV pipeline does not have precise road segmentation capabilities, it cannot conduct lane-level vehicle detection, failing to differentiate between parking and through vehicles. In such cases, observations provide link density for all vehicles on the link and we aggregate parking and through vehicles together to model density, which is already described in Equation (7) and (8). In this study, we integrate the aggregation into density estimation in the DNL model. Another general way is to estimate separate parking density and through traffic density in the DNL and using the matrix $\bm{I}_{i}$ to aggregate them. Another reason we aggregate them in the DNL process is that one through vehicle captured by a satellite image only represent the status at the timestamp of the image, and we are not sure if this vehicle will perform on-street parking after this timestamp or just left a parking spot before this timestamp. Aggregating through and parking vehicles into total density can overcome this information gap. The second scenario is about vehicle moving directions. Similar to scenario 1, if the CV pipeline cannot perform road segmentation and detect vehicle orientations, detected vehicles cannot be correctly matched to directed road segments, potentially leading to flow inconsistency on successive links. To address this, we aggregate two (even more) moving directions together to match regional total density. Another scenario is about vehicle class detection. When the CV pipeline fails to detect vehicle classes with high precision, we can simply aggregate density across vehicle classes.
This study extends the work of Liu et al. [2024] in both modeling and image data processing. This study incorporates curbside parking in dynamic network loading to explicitly model densities of both parking and moving vehicles, consistent with what satellite imagery captures, while Liu et al. [2024] did not differentiate moving and parking vehicles. Moreover, this study employs a class-specific vehicle detection algorithm to process satellite imagery data, improving the fidelity of density observations and enabling multi-class DODE, while Liu et al. [2024] used aggregated density for all vehicle classes. Note that the results of this study and Liu et al. [2024] cannot be compared directly due to the inherently underdetermined nature of DODE and different “ground truth” inputs.